# HateDay: Insights from a Global Hate Speech Dataset Representative of a Day on Twitter

### v1

@@ Abstract
To tackle the global challenge of online hate speech, a large body of research has developed detection models to flag hate speech in the sea of online content. Yet, due to systematic biases in evaluation datasets, detection performance in real-world settings remains unclear, let alone across geographies. To address this issue, we introduce HateDay, the first global hate speech dataset representative of social media settings, randomly sampled from all tweets posted on September 21, 2022 for eight languages and four English-speaking countries. Using HateDay, we show how the prevalence and composition of hate speech varies across languages and countries. We also find that evaluation on academic hate speech datasets overestimates real-world detection performance, which we find is very low, especially for non-European languages. We identify several factors explaining poor performance, including models’ inability to distinguish between hate and offensive speech, and the misalignment between academic target focus and real-world target prevalence. We finally argue that such low performance renders hate speech moderation with public detection models unfeasible, even in a human-in-the-loop setting which we find is prohibitively costly. Overall, we emphasize the need to evaluate future detection models from academia and platforms in real-world settings to address this global challenge.
@@ 1 Introduction
Social media users often encounter hate speech in their feeds UNESCO (2023), fueling fears of its impacts on offline violence Müller and Schwarz (2021). To counter this phenomenon, a mature body of research has developed automatic hate speech detection models Vidgen and Derczynski (2020). Yet, past work in this space suffers from three major shortcomings.
@@
Figure 1: HateDay consists of twelve annotated representative sets (N=20K each) randomly sampled from all tweets posted on September 21, 2022 in eight languages (Arabic, English, French, German, Indonesian, Portuguese, Spanish, and Turkish) and originating from four countries where English is the main language on Twitter (United States, India, Nigeria, Kenya).
First, the performance of hate speech detection systems in real-world settings remains largely unclear. On the one hand, detection models used in industry are not public and detection performance metrics are usually missing from platform transparency reports or sometimes misleading Giansiracusa (2021). On the other hand, evaluation of hate speech models in academic research has been done on biased datasets whose characteristics greatly differ from real-world settings, from class imbalance to topic distribution Nejadgholi and Kiritchenko (2020), raising concerns about an overestimation of the reported detection performance Arango et al. (2019); Wiegand et al. (2019).
@@
The second major limitation of academic research on automatic detection of hate speech lies in its limited linguistic scope, with most resources developed for English Poletto et al. (2021); Tonneau et al. (2024a). While such unequal resource allocation is likely to result in unequal performance across languages, the use of customized datasets for evaluation renders cross-lingual performance comparisons unfeasible.
@@
Finally, the focus on languages in the development of hate speech detection systems may hide performance differences between countries that share a common language Ghosh et al. (2021). While recent work has focused on developing datasets grounded in geographical location rather than language Maronikolakis et al. (2022); Castillo-lópez et al. (2023), cross-country performance differences remain largely unclear. In this context, it is necessary to develop multilingual evaluation sets that are representative of real-world settings and that provide a comparable setting across languages and countries with a common language.
@@
In this article, we aim to address these three shortcomings and provide an evaluation of state-of-the-art hate speech detection performance on real-world social media data across languages and countries. For this purpose, we introduce HateDay, a global Twitter dataset consisting of 240,000 tweets sampled from all tweets posted on September 21, 2022 and annotated for hate speech detection. Specifically, we draw random samples of 20,000 tweets each for eight popular languages on Twitter: Arabic, English, French, German, Indonesian, Portuguese, Spanish, and Turkish. In addition, we draw random samples of 20,000 tweets each for four countries whose major language on Twitter is English: the United States, India, Nigeria, and Kenya.
We show that the prevalence and composition of hate varies across languages and countries. We then use HateDay to evaluate public hate speech detection models across languages and countries. We find that hate speech detection performance is largely overestimated by evaluation on academic hate speech data from prior work and is very low in real-world social media settings, particularly for non-European languages. We identify several factors explaining such poor model performance including the inability of models to distinguish between hate and offensive speech, and the misalignment between academic target coverage and real-world target prevalence. Finally, in light of our performance results, we estimate the feasibility of hate speech moderation using publicly available models. We argue that automatic moderation alone is too error-prone to be desirable and show that human-in-the-loop moderation, whereby models flag potentially hateful content for human verification, proves prohibitively costly across languages and countries, despite its potential for improved accuracy. In sum, our contributions are:
@@
    HateDay, the first global hate speech dataset representative of social media settings [^1] Available at https://huggingface.co/datasets/manueltonneau/hateday, composed of 240,000 annotated tweets covering eight languages and four countries.
@@
    a comparison of hate speech prevalence and composition across languages and countries
@@
    an estimation of cross-lingual and geographic hate speech real-world detection performance, including a detailed analysis of pitfalls of current detection models
@@
    an estimation of the feasibility of hate speech moderation using publicly available models
@@ 2.1 Data Collection
We filter TwitterDay both at the language and country levels to enable a more structured analysis. At the language level, we focus on the eight most popular languages in the TwitterDay dataset for which there exist hate speech detection resources (details in §A.1), namely Arabic, English, French, German, Indonesian, Portuguese, Spanish and Turkish. To assess differences between countries with a common language, we also filter at the country-level, focusing on four countries for which English is the main language on Twitter, namely the United States, India, Nigeria and Kenya. For locating users in each country, we use the Google Geocoding API to infer user country location based on their self-provided profile location Hecht et al. (2011). We drop retweets to focus on original content and randomly sample 20,000 tweets for each of the eight retained languages and each of the four countries, ensuring that each language-specific and country-specific sample is representative of real-world Twitter settings in that particular language or country. The HateDay dataset corresponds to the combination of all the annotated random samples for each language and country, for a total of 240,000 annotated tweets.
@@ 2.2 Data Annotation
We recruit a team of 36 annotators, three per language or country. For languages that are spoken widely around the globe, such as English, Arabic or French, we maximize the diversity of annotator origins where possible (details in §B.1).
We follow a prescriptive approach to manage annotator subjectivity in our annotation task Rottger et al. (2022) by instructing annotators to strictly adhere to extensive annotation guidelines that describe our taxonomy of hate speech (provided in §B.2). Following prior work Davidson et al. (2017); Mathew et al. (2021), the annotation consists of assigning each tweet to one of three classes: (i) hateful, if it promotes violence or contains a direct attack, abuse or threat against an individual or a group based on the perceived belonging of a certain characteristic (e.g., gender, race), (ii) offensive, if it contains any form of non-acceptable language that is not hateful, including insults, threats, and posts containing profane language or swear words Zampieri et al. (2019), or (iii) neutral if it is neither hateful nor offensive. For tweets labeled as hateful, we also ask annotators to specify the groups being targeted. We provide examples of tweets for each class in Table 1.
Class Target Category Examples Origin
@@
For each language and country, we conduct a pilot annotation to train annotators on the task. Specifically, we have each annotator label 100 posts sampled from a collection of all annotated hate speech datasets Tonneau et al. (2024a) open-sourced on Hugging Face, a major open source platform. We repeat this task until the inter-annotator agreement, measured by Krippendorff’s $\alpha$, reaches 0.7. After the pilot, we task each annotator to label the random sample of 20,000 tweets in their respective language or country. Each tweet is labeled by three annotators, and the final label is determined by majority vote. To minimize the amount of hate that is wrongly categorized as not hate, all tweets without a majority vote or with at least one hateful label out of three are re-examined by all three annotators and a task expert to derive their final label. Across all languages and countries, the three annotators agreed on 93.4% of all labeled tweets, two out of three agreed in 6.4% of cases, and all three of them disagreed in 0.2% of cases.
@@ 2.3 Descriptive Statistics
We find that the prevalence of hate speech on Twitter on the day of analysis is very low, representing less than 2% of all posts for all considered languages and countries and with an average prevalence of 0.7% (Figure 2, left). In contrast, we observe that offensive content is much more prevalent than hate speech across all languages and countries, ranging from 2.5 times more prevalent for French to 19 times more prevalent for Portuguese. We also observe large differences in the prevalence of hate speech between languages and countries. For instance, 1.0% of German tweets are hateful whereas only 0.3% of Indonesian tweets are hateful. We observe similar gaps at the country level, with the share of hateful posts being much higher in India (1.7%) than in Kenya (0.2%) or Nigeria (0.3%).
@@
We find that the most common targets of hate in HateDay are political as well as racial and gender groups (Figure 2, right). We also find notable differences across languages and countries. For instance, political hate speech is prevalent in English, German, Indonesian, Portuguese, and Spanish, as well as in the U.S. national context, whereas it is less present in Turkish and Nigerian tweets, and not present at all in Arabic tweets. Also, religious hate speech, mostly in the form of Islamophobia, dominates in India whereas it is less prominent elsewhere and completely absent from the sample in the Kenyan case. Finally, we note that some forms of hate speech are unique to specific contexts, such as casteism in India.
@@ 3.1 Models
For each language and country, we identify and evaluate all hate speech detection models that are open-sourced on Hugging Face, and trained using supervised learning. In cases where there are more than five models, we limit our analysis to the five most downloaded models on Hugging Face at the time of analysis (August 2024). We provide the full list of open-source models that we evaluated in §D.1. Additionally, we include the Perspective API Lees et al. (2022), a widely used toxic language detection system also based on supervised learning.
@@
We use Aya23 8B Aryabumi et al. (2024) and Llama3.1 8B Dubey et al. (2024) for zero-shot learning. We do so because Aya is designed to be multilingual, and Llama3.1 is one of the most capable open-source decoder-only models available at the time of the analysis. We use smaller architectures due to computing constraints and provide the prompt for zero-shot learning in §D.2.
@@ 3.2 Evaluation
We evaluate models on the HateDay (HD) datasets to estimate real-world detection performance. Additionally, we compare the latter with the performance on two traditional evaluation methods, namely academic hate speech datasets and functional tests.
@@
We measure performance on academic hate speech datasets to understand how results from past work generalize to our more realistic HateDay datasets. We rely on open-sourced hate speech supersets combining all existing hate speech datasets for all eight languages of interest Tonneau et al. (2024a). Given that English hate speech datasets mostly originate from the US Tonneau et al. (2024a), we use English hate speech datasets both for evaluating English-centered and US-centered models. In the absence of supersets for India, Nigeria and Kenya, we survey all existing datasets for each country and combine them to build the supersets (details in §C.2). We restrict the evaluation to a 10% random sample of all academic hate speech datasets for each language and country to limit the computing resources required for inference (details in §C.1).
@@
We measure performance on functional tests to estimate models’ ability to handle known challenges in hate speech detection. We use HateCheck Röttger et al. (2021, 2022), a suite of functional tests for hate speech detection models, which covers six of eight languages in HateDay, missing Indonesian and Turkish.
We evaluate model performance on each dataset using average precision, which corresponds to the area under the precision-recall curve, and is well suited when class imbalance is high, as is the case in HateDay.
@@ 4.1 Detection Performance
Our most striking finding is that across the wide range of models, languages and countries considered in this study, the average precision on our representative HateDay data (HD) is just 9.4% on average, compared to 40% on the sample of public hate speech datasets (AD) and 87.2% on the functional tests (HC).
@@
We find that supervised learning consistently outperforms zero-shot learning across almost all combinations of language, country, and evaluation sets. Indeed, the best open-source model on HateDay has an average precision of 36.0% on average across all combinations, whereas Llama3.1 8B, which performs on par with Aya23 8B, scores 28.4% on average. Additionally, despite being originally developed to detect toxic rather than hateful content, we observe that the Perspective API (39.1% average precision on average) outperforms open-source hate speech detection models in all languages but German. Still, the dominance of Perspective is more limited at the country-level where it performs best in the US and Indian contexts but not in the Nigerian context, and only in the Kenyan context because there are no supervised models to compare to.
@@
Table 3: Model performance across countries and evaluation sets, as measured by average precision (%). There is no open-source model specifically for Kenya.
@@
We observe substantial performance differences in the real-world setting of HateDay across languages and countries. At the language-level, we find that the best average precision for European languages (23.1% on average), such as French, Spanish or German, is substantially higher than for non-European languages (9.3%), such as Arabic, Indonesian and Turkish. At the country-level, the average precision is by far highest on Indian tweets (42.9%) followed by Nigeria (30.9%), whereas the performance is rather low in both the U.S. (12.3%). and Kenyan (9.1%) contexts.
4.2 Performance Analysis
In the following, we conduct a more in-depth analysis into potential explanations for the poor hate speech detection performance as well as the cross-geographic performance gaps that we observe in real-world settings in §4.1.
@@ 4.2 Performance Analysis
As the average precision corresponds to the area under the precision-recall curve, the same average precision may correspond to very different patterns in precision and recall. We therefore inspect the precision-recall curves of the best-performing models on HateDay for each language and country. We find that, while precision or recall may be higher depending on the context, both values remain very low and there are no situations where both values are above 50%, with the highest F1-score at 47% for India (Figure 5 in the Appendix). This highlights the role of both false positives and negatives in the observed low performance. Next, we study the composition of these two types of errors.
Figure 3: Share of hateful, offensive and neutral content in the top 5% scored tweets (N=1,000) in HateDay for each language and country. We use the score distribution of the best performing model on HateDay (Tables 2 and 3). The x-axis corresponds to the descending rank and each bar shows the distribution of content in a window of 50 tweets.
@@
We find that offensive content constitutes a substantial share of the top of the score distribution — 32.3% in the top 50 (0.25%) scored tweets on average — for each language and country (Figure 3), thereby hurting precision. For Arabic, English, Portuguese and the U.S. and Kenyan national contexts, offensive content even appears more frequently than hate speech at the top of the score distribution. This problem is further aggravated by the fact that offensive content is more prevalent than hate speech on the day of the analysis (Figure 2). Indeed, we observe that as the prevalence ratio of offensive content to hate speech increases, the share of offensive content also gradually increases (Figure 7 in the Appendix).
@@
We further examine top-scored neutral tweets to identify other prominent features of false positives. We find that many such tweets contain mentions of hate speech (e.g., “wasn’t it the political scientist saying that Ukrainians are “subhumans”?”; “for them, women are like joints, they smoke them until they get high”). We also find several cases of statements that are not hateful but that would become hateful with a few changes of letters (e.g., “emigration is a catastrophy”; “this was noggers, man [^2] Portuguese original is: “isso foi noggers mano.”. “Noggers” means “disappointing” in Portuguese gamer slang.). Finally, we find that a substantial share of such tweets may be hateful but lack the context to fully determine their hatefulness. Such tweets often are replies and contain target community names (e.g., “@USER No, illegal migrants, they say it themselves”; “@USER We know you are gay”).
@@
Next, we analyze false negatives, defined as hateful examples that do not appear in the top 1% of the best model’s score distribution for each language and country. We first examine the targets of these tweets to determine how their representation in false negatives compares to their overall prevalence in hate speech for each specific language or country. We find that politics and gender are overrepresented in false negatives, respectively in 10/12 and 7/12 of the studied language and country-level cases. In parallel, religion and race are under-represented, respectively in 10/12 and 8/12 of cases. We also look at ambiguity, considering that a hateful example is ambiguous if not all three annotators labeled it as hateful. The share of ambiguous content is significantly higher in false negatives compared to the whole set of hateful examples across languages and countries (t(11) = 3.76; p-value < 0.05).
@@
Motivated by the overrepresentation of certain targets in false negatives, we inspect the role of target-level academic focus, as a proxy for the distribution of hate types seen by the model during training, on performance. We first examine differences at the target level between the attention each target group has received in academic work and its prevalence in HateDay across languages and countries. We use data from a recent survey which documents the share of all hate speech datasets by target category and language Yu et al. (2024). In the absence of data at the country level, we categorize all surveyed datasets in terms of the target focus for each country of interest (details in §C.2.2). For each language and country, we provide a comparison between the target-level shares of (i) all hate speech datasets and of (ii) all hate speech in HateDay in Figure 6 in the Appendix. Our most notable finding is that political hate speech is much more prevalent in HateDay in two-thirds of cases than it is studied in academic work on hate speech. To a lesser extent, we also find that gender-based hate is often more prevalent in HateDay than is studied while religion-based hate is less prevalent than is studied in existing hate speech datasets.
Finally, we explore the impact of the alignment between the target focus of academic work and the target prevalence in real-world social media data on performance. We estimate such alignment for each language and country by computing the cosine similarity between the vectors containing the share of all hate speech datasets by target focus and the share of all hate speech in HateDay by target. We find a strong and significant positive correlation between this alignment and detection performance (Pearson’s r: 0.76, p-value: 0.003). This indicates that the better hate speech detection resources reflect hate target coverage in real-world social media data for a given language or country, the higher the detection performance. In contrast, we do not find a significant correlation between the amount of annotated datapoints in the aforementioned supersets and performance on HateDay (Pearson’s r: -0.47, p-value: 0.127). This indicates that the amount of detection resources plays a smaller role in cross-lingual and geographic performance differences.

### v2

@@ Abstract
To address the global challenge of online hate speech, prior research has developed detection models to flag such content on social media. However, due to systematic biases in evaluation datasets, the real-world effectiveness of these models remains unclear, particularly across geographies. We introduce HateDay, the first global hate speech dataset representative of social media settings, constructed from a random sample of all tweets posted on September 21, 2022 and covering eight languages and four English-speaking countries. Using HateDay, we uncover substantial variation in the prevalence and composition of hate speech across languages and regions. We show that evaluations on academic datasets greatly overestimate real-world detection performance, which we find is very low, especially for non-European languages. Our analysis identifies key drivers of this gap, including models’ difficulty to distinguish hate from offensive speech and a mismatch between the target groups emphasized in academic datasets and those most frequently targeted in real-world settings. We argue that poor model performance makes public models ill-suited for automatic hate speech moderation and find that high moderation rates are only achievable with substantial human oversight. Our results underscore the need to evaluate detection systems on data that reflects the complexity and diversity of real-world social media.
@@ 1 Introduction
Social media users frequently encounter hate speech in their feeds UNESCO (2023), raising concerns about its potential to incite offline violence Müller and Schwarz (2021). In response, a substantial body of research has focused on developing automated hate speech detection models Vidgen and Derczynski (2020). However, prior work in this area suffers from three major limitations.
@@
Figure 1: HateDay consists of twelve annotated representative sets (N=20K each) randomly sampled from all tweets posted on September 21, 2022. The dataset covers eight languages (Arabic, English, French, German, Indonesian, Portuguese, Spanish, and Turkish) and four countries where English is the main language on Twitter (India, Kenya, Nigeria, and the United States).
First, the performance of hate speech detection systems in real-world social media settings remains largely unknown. Industry models are typically not publicly available, and platform transparency reports omit detailed performance metrics or present them in ways that can be misleading or difficult to interpret Giansiracusa (2021). Meanwhile, academic evaluations rely on biased datasets that diverge considerably from real-world distributions, especially regarding class imbalance and topic diversity Nejadgholi and Kiritchenko (2020). This raises concerns that reported performance may be substantially overestimated and not generalize well outside such controlled settings Arango et al. (2019); Wiegand et al. (2019).
@@
The second key limitation of academic research on automatic hate speech detection lies in its narrow language focus. Indeed, most available datasets and detection models have been developed for English Poletto et al. (2021); Tonneau et al. (2024a). While such unequal resource allocation is likely to result in unequal performance across languages, the widespread use of customized datasets for evaluation hinders meaningful cross-lingual comparisons.
@@
Finally, focusing on languages in hate speech detection can hide performance differences between countries sharing a language Ghosh et al. (2021). Although recent work develops datasets based on geography rather than language Maronikolakis et al. (2022); Castillo-lópez et al. (2023), cross-country performance variations remain unclear. Addressing this limitation requires evaluation datasets providing a comparable evaluation setting across countries with a common language.
@@
In this article, we address these three limitations by evaluating the performance of state-of-the-art hate speech detection models on real-world social media data across multiple languages and countries. We introduce HateDay, a global Twitter dataset which consists of 240,000 tweets, randomly sampled from all tweets posted on September 21, 2022 and annotated for hate speech (Figure 1, left). Specifically, we sample 20,000 tweets each for eight widely used languages on Twitter: Arabic, English, French, German, Indonesian, Portuguese, Spanish, and Turkish. We also sample 20,000 tweets each for four English-speaking countries with a sizable Twitter presence: India, Kenya, Nigeria, and the United States.
Using HateDay, we provide insights on real-world online hate (Figure 1, right). We show that the prevalence and composition of hate vary significantly across languages and countries. We then evaluate publicly available hate speech detection models across this global landscape. Our findings reveal that detection performance is substantially overestimated when assessed on standard academic datasets and is strikingly low in real-world social media settings—especially for non-European languages. We identify several factors contributing to this poor performance, including models’ limited ability to distinguish hate from offensive speech, and a mismatch between the target categories prevalent in academic datasets and those prevalent in real-world discourse.
Given these results, we examine the feasibility of using public models for hate speech moderation. We argue that automatic moderation—where posts flagged as hateful are moderated without human oversight—is currently too error-prone to be a viable approach. In contrast, we find that human-in-the-loop systems, in which flagged content is reviewed by moderators, can achieve higher moderation rates across languages and countries, albeit at the cost of substantial human review.
In sum, our contributions are:
@@
    HateDay, the first global hate speech dataset representative of social media settings, composed of 240,000 annotated tweets covering eight languages and four countries [^1] Available at https://huggingface.co/datasets/manueltonneau/hateday
@@
    a cross-lingual and cross-national comparison of hate prevalence and composition
@@
    an evaluation of real-world detection performance across languages and countries, along with a detailed analysis of the limitations of current detection models
@@
    an assessment of the feasibility of hate speech moderation using publicly available models, including a comparison of automatic and human-in-the-loop approaches
2 The HateDay Dataset
@@ 2.1 Data Collection
We filter TwitterDay both at the language and country levels. At the language level, we focus on the eight most popular languages in the TwitterDay dataset for which there exist hate speech detection resources (details in §A.1), namely Arabic, English, French, German, Indonesian, Portuguese, Spanish and Turkish. To assess differences between countries with a common language, we also filter at the country-level, focusing on four countries for which English is the main language on Twitter, namely India, Kenya, Nigeria, and the United States. We use the Google Geocoding API to infer user country location based on their self-provided profile location, knowing this has limitations Hecht et al. (2011). We drop retweets to focus on original content and randomly sample 20,000 tweets for each of the eight retained languages and each of the four countries, ensuring that each language-specific and country-specific sample is representative of real-world Twitter settings in that particular language or country. The HateDay dataset corresponds to the combination of all annotated random samples for each language and country, totaling 240,000 annotated tweets.
@@ 2.2 Data Annotation
We recruit a team of 36 annotators, three per language or country. For languages that are spoken widely around the globe, such as English, Arabic or French, we maximize the diversity of annotator origins as much as possible (details in §B.1).
We follow a prescriptive approach to manage annotator subjectivity in our annotation task Röttger et al. (2022b) by instructing annotators to strictly adhere to extensive annotation guidelines that describe our taxonomy of hate speech (provided in §B.2). Following prior work Davidson et al. (2017); Mathew et al. (2021), the annotation consists of assigning each tweet to one of three classes: (i) hateful, if it promotes violence or contains a direct attack, abuse or threat against an individual or a group based on the perceived belonging of a certain characteristic (e.g., gender, race), (ii) offensive, if it contains potentially objectionable language that is not hateful, including insults, threats, and posts containing profanity or swear words Zampieri et al. (2019), or (iii) neutral if it is neither hateful nor offensive. For tweets labeled as hateful, we also ask annotators to specify the groups being targeted as a free-text label. We provide examples of tweets for each class in Table 1.
Class Target Category Examples Split
@@
For each language and country, we conduct a pilot annotation phase to train annotators on the task. Specifically, we have each annotator label 100 posts, sampled from a collection of all annotated hate speech datasets Tonneau et al. (2024a), which have been open-sourced on Hugging Face. We repeat this task until the inter-annotator agreement, measured by Krippendorff’s $\alpha$, reaches a threshold of 0.7. After completing the pilot, we then task each annotator with labeling the random sample of 20,000 tweets in their respective language or country. Each tweet is labeled by three annotators, and the final label is determined by a simple majority vote. Across all languages and countries, the three annotators agreed on 93.4% of all labeled tweets; two out of three agreed in a further 6.4% of cases, and all three disagreed in only 0.2% of cases.
@@ 2.3 Descriptive Statistics
We find that the prevalence of hate speech on Twitter on the day of analysis is very low, representing less than 2% of all posts across languages and countries, with an average prevalence of 0.7% (Figure 2, left). In contrast, offensive content is substantially more prevalent than hate speech across all considered languages and countries, ranging from 2.5 times more prevalent for French to 19 times more prevalent for Portuguese. We also find notable differences in the prevalence of hate speech between languages and countries. For instance, 1.0% of German tweets are hateful whereas only 0.3% of Indonesian tweets are hateful. We observe similar gaps at the country level, with the share of hateful posts being much higher in India—1.7%—compared to 0.2% in Kenya and 0.3% in Nigeria.
@@
We find that the most common targets of hate in HateDay are political as well as racial and gender groups (Figure 2, right). We also find notable differences across languages and countries. For instance, political hate speech is prevalent in English, German, Indonesian, Portuguese, and Spanish, as well as in the US national context, representing up to 66% of all hate in the German context. In contrast, it is less present in Turkish and Nigerian tweets, and not present at all in Arabic tweets. Also, religious hate speech represents 41% of all hate in India, mostly in the form of Islamophobia, whereas it is less prominent elsewhere and completely absent from the Kenyan sample. Finally, we note that some forms of hate speech are unique to specific contexts, such as casteism in India.
@@ 3.1 Models
For each language and country, we identify and evaluate all hate speech detection models that are open-sourced on Hugging Face, and trained using supervised learning. In cases where there are more than five models, we limit our analysis to the five most downloaded models on Hugging Face at the time of analysis (August 2024). We provide the full list of open-source models that we evaluated in §D.1. Additionally, we include the Perspective API Lees et al. (2022), a widely used toxic language detection system also based on supervised learning. Specifically, we use the API’s “Identity Attack” attribute, as it most closely aligns with our definition of hate speech.
@@
We use Aya23 8B Aryabumi et al. (2024) and Llama3.1 8B Dubey et al. (2024) for zero-shot learning. We do so because Aya is designed to be multilingual, and Llama3.1 is one of the most capable open-source models available at the time of our analysis. We use small model sizes due to compute constraints. For the zero-shot prompt, see §D.2.
@@ 3.2 Evaluation
We evaluate models on the HateDay (HD) datasets to estimate real-world detection performance. For comparison, we also test models in two traditional evaluation settings, namely on academic hate speech datasets and functional tests.
@@
We measure performance on academic hate speech datasets to understand how results from past work generalize to more realistic settings. We rely on supersets combining all existing hate speech datasets for all eight languages of interest Tonneau et al. (2024a). Given that English hate speech datasets mostly originate from the US Tonneau et al. (2024a), we use English hate speech datasets both for evaluating English- and US-centered models. In the absence of supersets for India, Nigeria and Kenya, we survey all existing datasets for each country and combine them to build the supersets (details in §C.2). We restrict the evaluation to a 10% random sample of all academic datasets for each language and country to limit inference costs (details in §C.1).
@@
We measure performance on functional tests to estimate the ability of models to handle known challenges in hate speech detection. We use HateCheck Röttger et al. (2021, 2022a), a suite of functional tests for hate speech detection models, which covers six of eight languages in HateDay, missing Indonesian and Turkish.
Evaluation metric
We evaluate model performance using average precision, which corresponds to the area under the precision-recall curve, and is well suited when class imbalance is high.
@@ 4.1 Detection Performance
Our most striking finding is that across the wide range of models, languages, and countries considered, detection performance is much lower on our representative HateDay dataset (HD) than on academic hate speech datasets (AD) or functional tests (HC). Indeed, average precision is just 9.4% on HD, compared to 40% on AD and 87.2% on HC.
@@
We find that supervised learning consistently outperforms zero-shot learning across almost all combinations of language, country, and evaluation sets. Indeed, the best open-source model on HateDay has an average precision of 41.1% on average across all combinations of language or country and dataset, whereas Aya23 8B, which performs on par with Llama3.1 8B, scores 32.4% on average. Additionally, despite being originally developed to detect toxic rather than hateful content, we observe that the Perspective API (44.3% average precision on average) outperforms open-source hate speech detection models in all languages except German. At the country-level, open-source models perform best in Nigeria, but Perspective API has higher performance in the US and India. We did not find open-source supervised models to compare Perpective API to in the Kenyan context.
@@
Table 3: Model performance across countries and evaluation sets, as measured by average precision (%). There are no open-source models specifically for Kenya.
@@
We observe substantial performance differences in the real-world setting of HateDay across languages and countries. At the language level, average precision is higher for European languages—English, French, Spanish, Portuguese, and German—averaging 23.1%—than for non-European languages—Arabic, Indonesian, and Turkish—at 9.3% on average. At the country level, performance is highest on Indian tweets (42.9%), followed by Nigeria (30.9%), and lower in both the US (12.3%) and Kenya (9.1%).
4.2 Reasons for Low Performance
In the following, we conduct a more in-depth analysis into potential explanations for the poor hate speech detection performance as well as the cross-geographic performance gaps we observed in real-world settings in §4.1.
@@ 4.2 Reasons for Low Performance
As the average precision corresponds to the area under the precision-recall curve, the same average precision may correspond to very different patterns in precision and recall. We therefore inspect the precision-recall curves of the best-performing models on HateDay for each language and country (Figure 5 in the Appendix). We find that, while precision or recall may be higher depending on the context, both values remain very low and there are no situations where both values are above 50%, with the highest F1-score at 47% for India. This highlights the role of both false positives and negatives in the observed low performance. Next, we study the composition of these two types of errors.
Figure 3: Share of hateful, offensive and neutral content in the top 5% scored tweets (N=1,000) in HateDay for each language and country. We use the hatefulness score distribution of the best performing model on HateDay (Tables 2 and 3). The x-axis corresponds to the descending rank and each bar shows the distribution of content in a window of 50 tweets.
@@
We find that offensive content constitutes a substantial share of the top of the hatefulness score distribution—32.3% of the top 50 (0.25%) scored tweets on average—for each language and country (Figure 3), thereby significantly hurting precision. For Arabic, English, Portuguese, and the US and Kenyan national contexts, offensive content even appears more frequently than hate speech at the very top of the score distribution. This problem is further aggravated by the fact that offensive content is more prevalent than hate speech (Figure 2). Indeed, we observe that as the prevalence ratio of offensive content to hate speech steadily increases, the share of offensive content at the top of the score distribution also gradually increases (Figure 7 in the Appendix), exacerbating the negative impact on precision.
@@
We further examine top-scored neutral tweets to identify other prominent features of false positives beyond offensiveness. We find that many such tweets contain mentions of hate speech (e.g., “was it the political scientist saying that Ukrainians are ‘subhumans’?”). We also find several cases of statements that are not hateful but would be with a few changes of letters (e.g., “emigration is a catastrophy”). Finally, we find that a substantial share of such tweets, usually replies, may be hateful but lack the context to conclude (e.g., “@USER illegal migrants, they say it themselves”).
@@
Next, we analyze false negatives, defined as hateful examples missing from the top 1% of the best model’s score distribution for each language and country. We first examine the targets of these tweets to compare their representation in false negatives to their overall prevalence in hate speech for each language or country. We find that politics and gender are overrepresented in false negatives, respectively in 10/12 and 7/12 of cases. Conversely, religion and race are underrepresented in 10/12 and 8/12 of cases. We also consider ambiguity, defining a hateful example as ambiguous if not all three annotators labeled it as hateful. We find that the share of ambiguous content is significantly higher in false negatives than in all hateful examples (t(11) = 3.76; $p<0.05$).
@@
Motivated by the overrepresentation of certain targets in false negatives, we inspect the role of target-level academic focus, as a proxy for the distribution of hate targets seen by the model during training, on performance. We first examine differences at the target level between the prevalence of each target in academic datasets and its prevalence in HateDay across languages and countries. We use data from a recent survey which documents the share of all hate speech datasets by target category and language Yu et al. (2024). In the absence of data at the country level, we categorize all surveyed datasets in terms of the target focus for each country of interest (details in §C.2.2). For each language and country, we provide a comparison between the target-level shares of (i) all hate speech datasets and of (ii) all hate speech in HateDay in Figure 6 in the Appendix. Our most notable finding is that political hate speech is much more prevalent in HateDay for two-thirds of the 12 languages or countries than in academic datasets. To a lesser extent, we also find that gender-based hate is often more prevalent in HateDay than is studied while religion-based hate is less prevalent than is studied in existing hate speech datasets.
Target alignment and performance
Finally, we explore the impact of the alignment between the target focus of academic work and the target prevalence in real-world social media data on performance. We estimate such alignment for each language and country by computing the cosine similarity between the vectors containing the share of all hate speech datasets by target focus and the share of all hate speech in HateDay by target. We find a strong and significant positive correlation between this alignment and detection performance (Pearson’s r: 0.76, $p=0.003$). This indicates that the better hate speech detection resources reflect hate target coverage in real-world social media data for a given language or country, the higher the detection performance. In contrast, we do not find a significant correlation between the amount of annotated datapoints in the aforementioned supersets and performance on HateDay (Pearson’s r: -0.47, $p=0.127$). This indicates that the amount of detection resources may play a smaller role in cross-lingual and geographic performance differences compared to the alignment between academic target focus and real-world target prevalence.